# A Momentum Balance Correction to the Non-Conservative One-Fluid Formulation in Boiling Flows using Volume-of-Fluid

### v1

@@ 4.1 One-dimensional Evaporating Liquid Surface
A numerical test following the one-dimensional liquid surface configuration described in Section 3.5 shows the recovery of the correct pressure and velocity jumps across $\Omega_{\Gamma}$ in the one-fluid formulation, regardless of the mesh resolution, as seen in Figure 5. The computational domain is defined with the wall at $x=0$, the liquid-gas interface initially at $x=25$ mm and an open boundary at $x=150$ mm where the reference pressure is set at 0 Pa. The mass flux is set to $\dot{m}^{\prime}=10$ kg/(m2s), resulting in $p_{l}=0.8$ Pa and $u_{g}=0.08$ m/s. Various mesh resolutions are tested, which are given in Table 1 (i.e., L1, L2, L3 and L4). Although the error in the momentum balance introduced by the one-fluid velocity in the viscous term is negligible (i.e., well below 1% for the tested grid sizes and fluid properties), the results in this section have been obtained by neglecting viscous effects in the momentum equation.
@@
The errors introduced by the one-fluid formulation without corrections (i.e., $\mathbf{f}_{NC}=0$, $\mathbf{f}_{\dot{m}^{\prime}}=0$ and $\mathbf{u}^{*}=\mathbf{u}^{n}$) are visualized in Figure 6 by comparing the results obtained with L3. As expected, a pressure field is obtained that corrects the predicted velocity to satisfy $\nabla\cdot\mathbf{u}^{n+1}$. Therefore, the velocity jump is the same with and without corrections (see Figure 6(a)). However, the pressure jump is overestimated by about 35% and varies over time as the interface moves across cells, as seen in Figures 6(b) and 6(c). Further, the figure also analyzes the individual effect of each correction. For this particular configuration, the momentum imbalance caused by the convective term does not vary much over time as the interface crosses grid cells; thus, the addition of $\mathbf{f}_{\dot{m}^{\prime}}$ and $\mathbf{f}_{NC}$ mainly correct the pressure offset necessary to obtain the exact $\Delta p$. In contrast, the Stefan flow shift is responsible for the elimination of the large temporal pressure oscillations that take some time to decrease before re-emerging once a new cell is suddenly identified as belonging to $\Omega_{\Gamma}$.
@@ 4.2 Static Evaporating Droplet
The flow solver is tested for a static droplet evaporating under constant $\dot{m}^{\prime}$ to analyze the physical consistency of the solution and stability of the proposed method. An analytical solution to a simplified version of this problem can be obtained by decoupling the velocity calculation from the pressure. Under spherical symmetry, the radial velocity $u_{r}$ is obtained from the continuity equation in spherical coordinates, resulting in
@@
where $u_{G}=\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})$. Next, pressure is calculated by substituting $u_{r}$ into the momentum equation, also in spherical coordinates. Assuming that $p_{r\rightarrow\infty}=0$, the pressure is given by
$p(r,t)=\begin{cases}p_{G}+\sigma\kappa+(\dot{m}^{\prime})^{2}(\rho_{G}^{-1}-\rho_{L}^{-1})+4\frac{\mu_{G}}{R(t)}u_{G}&\text{if $r<R(t)$}\\ -\frac{1}{2}\rho_{G}u_{G}^{2}\bigg{(}\frac{R(t)}{r}\bigg{)}^{4}&\text{if $r\geq R(t)$}\\ \end{cases}$ (40)
with $p_{G}=-\frac{1}{2}\rho_{G}u_{G}^{2}$. In Eqs. (39) and (40), the droplet radius is given as a function of time by $R(t)=R_{0}-\frac{\dot{m}^{\prime}}{\rho_{L}}t$ where $R_{0}$ is the initial droplet radius. Note the pressure jump due to the normal viscous stress is included despite being negligible compared to the other terms.
@@
A range of $\dot{m}^{\prime}$ and $\sigma$ values are considered (see Table 2) to test the accuracy of the one-fluid corrections and the interplay with surface tension. A non-dimensional time is defined as $t^{*}=t/t_{c}$, where the characteristic time is given by $t_{c}=\rho_{L}d_{0}/\dot{m}^{\prime}$. Note the scaling with the interface regression velocity $\mathbf{u}_{\Gamma}$ for a static droplet.
Case $\dot{m}^{\prime}$ [kg/(m2s)] $\sigma$ [mN/m]
@@
Table 2: Mass flux per unit area, $\dot{m}^{\prime}$, and surface tension coefficient, $\sigma$, for the different configurations of the static evaporating droplet.
@@ 4.2.1 Flow Field
The impact of not considering the flow solver corrections is seen in Figure 7(c) using L3. As expected from the discussion in Section 4.1, a bigger pressure jump is introduced without the corrections. Compared to the one-dimensional results shown in Figure 6, the pressure field is incorrect in both phases without the flow solver corrections. In particular, a larger pressure inside the droplet is observed, while the pressure in the gas phase is substantially lower than the analytical solution.
@@
All meshes capture the volume of the droplet accurately and maintain the spherical shape within resolution limits. However, a tendency to underpredict the droplet’s volume is observed in Figure 8(a), especially for $t^{*}>0.1$. This affects the pressure inside the droplet shown in Figure 8(b) as it shrinks and the pressure jump due to surface tension becomes more important, i.e., a larger curvature is predicted due to the smaller volume. A formal error analysis is presented in Section 4.2.2.
The evolution of the droplet shape is shown in Figure 9 for case 3 with mesh L2 together with the magnitude of $\mathbf{u}$, plotted on the $xy$ plane at $z=0$. The droplet remains spherical despite the spurious currents that appear around the interface, which deteriorate over time as the droplet resolution decreases. These affect the Stefan flow in the gas phase and some degree of non-symmetry is seen. 8 c/d has been observed to be the limit where shape deterioration becomes evident, spurious currents rise considerably and numerical oscillations in, e.g., the internal droplet pressure, significantly increase (see L1 in Figure 8(b)).
@@
Figure 9: Temporal evolution of the droplet represented by the iso-surface $C=0.5$ for case 3 with mesh L2. The velocity magnitude resulting from the Stefan flow is shown on the $xy$ plane across the center of the droplet. The length scale in each snapshot remains constant.
@@ 4.2.2 Numerical Stability and Error Analysis
The extension of the one-fluid formulation to two-phase flows undergoing phase change introduces numerical errors beyond the physical consistency of the results. The extensive use of non-smooth geometrical information (i.e., $\kappa$ and $\mathbf{n}_{\Gamma}$) from the interface reconstruction in the calculation of $\mathbf{f}_{\sigma}$, $\mathbf{f}_{NC}$ or the interface shift in Eq. (9) to account for phase change in the VOF advection generate spurious currents around the interface that further deteriorate $\mathbf{n}_{\Gamma}$ and $\kappa$. Moreover, these numerical oscillations easily affect the extrapolation of $\hat{p}$ and the stability of the flow solver. Partially affected by the low viscosities considered here, these issues are evident in the context of a static evaporating droplet where no convective flow exists and the velocities are solely a result of the Stefan flow and spurious currents.
@@
This motivates the use of PPIC to improve the calculation of $\kappa$, which is a more sensitive parameter than $\mathbf{n}_{\Gamma}$ under geometrical deterioration. In fact, from the cases outlined in Table 2, a direct benefit of using PPIC vs. PLIC is observed in achieving a smoother flow around the interface and stabilizing the solution throughout the entire droplet evaporation process (within resolution limits) for most grid sizes. With PLIC, case 1 eventually becomes unstable as spurious currents grow unbounded during many numerical integration steps. That is, as $\dot{m}^{\prime}$ decreases, the spurious currents become of the order of the Stefan flow and dominate the dynamics in $\Omega_{\Gamma}$. The use of energy-preserving schemes (i.e., balancing surface and kinetic energy) has been shown to guarantee the stability of the system [55] and may be pursued in future work.
@@
Error metrics in terms of convergence with grid refinement are evaluated for the droplet’s internal pressure and volume at $t^{*}=0.08$, as shown in Figures 10(a) and 10(b), respectively. Here, the mesh resolutions from Table 1 are also used. At this $t^{*}$, a large number of numerical integration steps have been taken and the volume has been reduced to $V/V_{0}=0.592$. Two scenarios are considered by fixing $\sigma=1$ mN/m and varying $\dot{m}^{\prime}$ (i.e., cases 1 and 3) and, then, by fixing $\dot{m}^{\prime}=10$ kg/(m2s) and varying $\sigma$ (i.e., cases 2 and 3). The relative error at $t^{*}=0.08$ is evaluated as $E=|\phi_{n}-\phi_{e}|/\phi_{e}$, where $\phi_{n}$ and $\phi_{e}$ are the value of any variable $\phi$ obtained, respectively, from the numerical solution and the exact (analytical) solution.
@@
Figure 10: Grid convergence of the internal droplet pressure and droplet volume at $t^{*}=0.08$ s in the static evaporating droplet. (a) pressure; (b) volume; (c) volume convergence for case 3 under the effect of the phase-change CFL constraint given by $\Delta\text{d}_{\text{max}}$.
@@
The convergence of the internal pressure error is nearly second order. This behavior is expected for a static droplet with $\dot{m}^{\prime}=0$ given the modeling of $\mathbf{f}_{\sigma}$ in the context of the CSF approach and shows that the addition of $\mathbf{f}_{\dot{m}^{\prime}}$ and $\mathbf{f}_{NC}$ behaves similarly. However, larger $\dot{m}^{\prime}$ values (cases 2 and 3) result in convergence rates varying between first and second order depending on the range of mesh sizes considered. Specifically, the convergence rate drops closer to first order when $\sigma\rightarrow 0$ (case 2). Assuming that $\mathbf{f}_{\dot{m}^{\prime}}$ behaves similar to $\mathbf{f}_{\sigma}$, it means the discretization of $\mathbf{f}_{NC}$ is first-order at best. This is a result of the stability (i.e., source term smoothness) trade-off between a CSF-like discretization and the approach described in Section 3.3 using staggered volume fractions to obtain the interfacial terms in $\mathbf{f}_{NC}$.
@@
Despite capturing the correct volume decrease and maintaining a spherical shape, grid convergence of the droplet’s volume is hardly achieved with the plane shift strategy along the normal direction described in Section 2.2. This is a clear consequence of the lack of smoothness in $\mathbf{n}_{\Gamma}$. For instance, Figure 10(b) shows that the method benefits from the smoother $\mathbf{n}_{\Gamma}$ distribution when the spurious currents in $\Omega_{\Gamma}$ are less, such as when $\dot{m}^{\prime}\rightarrow 0$ (case 1) or $\sigma\rightarrow 0$ (case 2). However, since grid refinement demands a larger number of time steps to reach the same $t^{*}=0.08$, the accumulation of geometrical errors during plane shifts prevents a clear converging trend. Nonetheless, all mesh resolutions from Table 1 display volume errors around 1% or less, in line with other works using different methodologies to handle phase change in VOF methods [11, 35].
@@
Note that this validation test imposes a restrictive CFL condition based on a maximum allowed interface shift due to phase change in a given time step of $(\Delta\text{d}_{\text{max}}/\Delta x)=0.001$. However, no significant improvement in the volume errors is observed by varying $(\Delta\text{d}_{\text{max}}/\Delta x)$ or $\theta_{3}$ from $0.04$ to $0.001$ (see C for the definition of $\theta_{3}$). Figure 10(c) shows the volume errors for case 3 and, similar to Figure 10(b), grid convergence is not achieved. Except for mesh L1, volume errors are almost identical regardless of $\Delta\text{d}_{\text{max}}$. Given these observations, a maximum plane shift around $(\Delta\text{d}_{\text{max}}/\Delta x)=0.01$ is more reasonable to avoid reducing the time step significantly if its calculation is limited by the evaporation of the interface.
The findings in this section highlight some limitations of the proposed treatment of phase change in the geometric advection of the volume fraction and the discretization of the various corrective source terms added in the momentum equation. Beyond the reported error trends, it has been found that reducing the time step delays the onset of numerical instabilities because of two main reasons: (a) smaller plane shifts in the phase change correction of $C$ deteriorate less the geometry; and (b) the errors in the temporal extrapolation of $\hat{p}$, $\hat{\mathbf{f}}_{\sigma}$ and $\hat{\mathbf{f}}_{\dot{m}^{\prime}}$ are reduced. Thus, spurious currents are smaller and geometrical calculations do not degrade as fast.

### v2

@@ 4.1 One-dimensional Evaporating Liquid Surface
A numerical test following the one-dimensional liquid surface configuration described in Section 3.5 shows the recovery of the correct pressure and velocity jumps across $\Omega_{\Gamma}$ in the one-fluid formulation, regardless of the mesh resolution, as seen in Figure 5. The computational domain is defined with the wall at $x=0$, the liquid-gas interface initially at $x=25$ mm and an open boundary at $x=150$ mm where the reference pressure is set at 0 Pa. The mass flux is set to $\dot{m}^{\prime\prime}=10$ kg/(m2s), resulting in $p_{l}=0.8$ Pa and $u_{g}=0.08$ m/s. Various mesh resolutions are tested, which are given in Table 1 (i.e., L1, L2, L3 and L4). Although the error in the momentum balance introduced by the one-fluid velocity in the viscous term is negligible (i.e., well below 1% for the tested grid sizes and fluid properties), the results in this section have been obtained by neglecting viscous effects in the one-fluid momentum equation.
@@
The errors introduced by the one-fluid formulation without corrections (i.e., $\mathbf{f}_{NC}=0$, $\mathbf{f}_{\dot{m}^{\prime\prime}}=0$ and $\mathbf{u}^{*}=\mathbf{u}^{n}$) are visualized in Figure 6 by comparing the results obtained with L3. Further, the figure also analyzes the individual effect of each correction. As expected, a pressure field is obtained that corrects the predicted velocity to satisfy $\nabla\cdot\mathbf{u}^{n+1}$. Therefore, the velocity jump is the same with and without corrections (see Figure 6(a)). However, the pressures are different. For example, the pressure difference $p^{*}$ between the pressures $p^{n+1}_{\text{FSM}}$ (W/o corrections) and $p^{n+1}$ (Stefan shift only), as defined in Section 3.4, can be as large as 1 Pa or 25% more than the actual pressure jump. Moreover, $p^{*}$ varies over time as the interface moves across cells, as seen in Figures 6(b) and 6(c). For this particular configuration, the momentum imbalance caused by the convective term does not vary much over time as the interface crosses grid cells; thus, the addition of $\mathbf{f}_{\dot{m}^{\prime\prime}}$ and $\mathbf{f}_{NC}$ mainly correct the pressure offset necessary to obtain the exact $\Delta p$ with $p^{n+1}$. In this case, the pressure jump is underestimated by more than 12.5% without the additional body forces. In contrast, the Stefan flow shift is responsible for the elimination of the large temporal pressure oscillations that take some time to decrease before re-emerging once a new cell is suddenly identified as belonging to $\Omega_{\Gamma}$. That is, $\Delta p$ spikes as the one-fluid suddenly accelerates due to the volumetric expansion caused by phase change when $\Omega_{\Gamma}$ moves across cells.
@@ 4.2 Static Evaporating Droplet
The flow solver is tested for a static droplet evaporating under constant $\dot{m}^{\prime\prime}$ to analyze the physical consistency of the solution and stability of the proposed method. An analytical solution to a simplified version of this problem can be obtained by decoupling the velocity calculation from the pressure. Under spherical symmetry, the radial velocity $u_{r}$ is obtained from the continuity equation in spherical coordinates, resulting in
@@
where $u_{G}=\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})$. Next, the pressure is calculated by substituting $u_{r}$ into the momentum equation, also in spherical coordinates. Assuming that $p_{r\rightarrow\infty}=0$ and a quasi-steady state evolution of the interface location, the pressure is given by
$p(r,t)=\begin{cases}p_{G}+\sigma\kappa+(\dot{m}^{\prime\prime})^{2}(\rho_{G}^{-1}-\rho_{L}^{-1})+4\frac{\mu_{G}}{R(t)}u_{G}&\text{if $r<R(t)$}\\ -\frac{1}{2}\rho_{G}u_{G}^{2}\bigg{(}\frac{R(t)}{r}\bigg{)}^{4}&\text{if $r\geq R(t)$}\\ \end{cases}$ (40)
with $p_{G}=-\frac{1}{2}\rho_{G}u_{G}^{2}$. In Eqs. (39) and (40), the droplet radius is given as a function of time by $R(t)=R_{0}-\frac{\dot{m}^{\prime\prime}}{\rho_{L}}t$ where $R_{0}$ is the initial radius. Note the pressure jump due to the normal viscous stress is included despite being negligible compared to the other terms.
@@
A range of $\dot{m}^{\prime\prime}$ and $\sigma$ values are considered (see Table 2) to test the accuracy of the one-fluid corrections and the interplay with surface tension. A non-dimensional time is defined as $t^{*}=t/t_{c}$, where the characteristic time is given by $t_{c}=\rho_{L}d_{0}/\dot{m}^{\prime\prime}$. Note the scaling with the interface regression velocity $\mathbf{u}_{\Gamma}$ for a static droplet.
Case $\dot{m}^{\prime\prime}$ [kg/(m2s)] $\sigma$ [mN/m]
@@
Table 2: Mass flux per unit area, $\dot{m}^{\prime\prime}$, and surface tension coefficient, $\sigma$, for the different configurations of the static evaporating droplet.
@@ 4.2.1 Flow Field
The impact of not considering the flow solver corrections is seen in Figure 7(c) using L3. As expected from the discussion in Section 4.1, a bigger pressure jump is introduced without the corrections, while the correct $\Delta p$ is captured with the corrections. Compared to the one-dimensional results shown in Figure 6, the pressure field is unphysical in both cases. In particular, the pressure field in the gas phase without corrections may be more consistent with the unsteady analytical solution but overestimates $\Delta p$, while the pressure field with corrections misses the unsteady effect of the shift in Stefan flow but reproduces the quasi-steady analytical solution and imposes the exact $\Delta p$.
@@
All meshes capture the volume of the droplet accurately and maintain the spherical shape within resolution limits. A similar behavior is shown in [13]. However, a tendency to underpredict the droplet’s volume is observed in Figure 8(a), especially for $t^{*}>0.1$. This affects the pressure inside the droplet shown in Figure 8(b) as it shrinks and the pressure jump due to surface tension becomes more important, i.e., a larger curvature is predicted due to the smaller volume. A formal error analysis is presented in Section 4.2.2.
The evolution of the droplet shape is shown in Figure 9 for case 3 with mesh L3 together with the magnitude of $\mathbf{u}$, plotted on the $xy$ plane at $z=0$. The droplet remains nearly spherical despite the spurious currents that appear around the interface, which deteriorate over time as the droplet resolution decreases. The sphericity of the droplet is shown in Figure 11. These affect the Stefan flow in the gas phase and some degree of non-symmetry is seen. 8 c/d has been observed to be the limit where shape deterioration becomes evident, spurious currents rise considerably and numerical oscillations in, e.g., the internal droplet pressure, significantly increase (see L1 in Figure 8(b)).
@@
Figure 9: Temporal evolution of the droplet represented by the iso-surface $C=0.5$ for case 3 with mesh L3. The velocity magnitude resulting from the Stefan flow is shown on the $xy$ plane across the center of the droplet. The length scale in each snapshot remains constant.
@@ 4.2.2 Numerical Stability and Error Analysis
The extension of the one-fluid formulation to two-phase flows undergoing phase change introduces numerical errors beyond the physical consistency of the results. The extensive use of non-smooth geometrical information (i.e., $\kappa$ and $\mathbf{n}_{\Gamma}$) from the interface reconstruction in the calculation of $\mathbf{f}_{\sigma}$, $\mathbf{f}_{NC}$ or the interface shift in Eq. (9) to account for phase change in the VOF advection generate spurious currents around the interface that further deteriorate $\mathbf{n}_{\Gamma}$ and $\kappa$. Moreover, these numerical oscillations easily affect the extrapolation of $\hat{p}$ and the stability of the flow solver. Partially affected by the low viscosities considered here, i.e., less dissipation, these issues intrinsic to geometric VOF frameworks are evident in the context of a static evaporating droplet where no convective flow exists and the velocities are solely a result of the Stefan flow and spurious currents.
@@
This motivates the use of PPIC to improve the calculation of $\kappa$, which is a more sensitive parameter than $\mathbf{n}_{\Gamma}$ under geometrical deterioration. In fact, from the cases outlined in Table 2, a direct benefit of using PPIC vs. PLIC is observed in achieving a smoother flow around the interface and stabilizing the solution throughout the entire droplet evaporation process (within resolution limits) for most grid sizes. With PLIC, case 1 eventually becomes unstable as spurious currents grow unbounded. That is, as $\dot{m}^{\prime\prime}$ decreases, the spurious currents become of the order of the Stefan flow and dominate the dynamics in $\Omega_{\Gamma}$. The use of energy-preserving schemes (i.e., balancing surface and kinetic energy) has been shown to guarantee the stability of the system [57] and may be pursued in future work.
@@
Error metrics in terms of convergence with grid refinement are evaluated for the droplet’s internal pressure and volume at $t^{*}=0.08$, as shown in Figures 10(a) and 10(b), respectively. Here, the mesh resolutions from Table 1 are also used. At this $t^{*}$, a large number of numerical integration steps have been taken and the volume has been reduced to $V/V_{0}=0.592$. Two scenarios are considered by fixing $\sigma=1$ mN/m and varying $\dot{m}^{\prime\prime}$ (i.e., cases 1 and 3) and, then, by fixing $\dot{m}^{\prime\prime}=10$ kg/(m2s) and varying $\sigma$ (i.e., cases 2 and 3). The relative error at $t^{*}=0.08$ is evaluated as $E=|\phi_{n}-\phi_{e}|/\phi_{e}$, where $\phi_{n}$ and $\phi_{e}$ are the value of any variable $\phi$ obtained, respectively, from the numerical solution and the exact (analytical) solution.
@@
Figure 10: Grid convergence of the internal droplet pressure and droplet volume at $t^{*}=0.08$ s in the static evaporating droplet. (a) pressure; (b) volume; (c) volume convergence for case 3 under the effect of the phase-change CFL constraint given by $\Delta\text{d}_{\text{max}}$. Case 3* corresponds to case 3 but without solving the momentum equation.
@@
The convergence of the internal pressure error is nearly second order. This behavior is expected for a static droplet with $\dot{m}^{\prime\prime}=0$ given the modeling of $\mathbf{f}_{\sigma}$ in the context of the CSF approach and shows that the addition of $\mathbf{f}_{\dot{m}^{\prime\prime}}$ and $\mathbf{f}_{NC}$ behaves similarly. However, larger $\dot{m}^{\prime\prime}$ values (cases 2 and 3) result in convergence rates varying between first and second order depending on the range of mesh sizes considered. Specifically, the convergence rate drops closer to first order when $\sigma\rightarrow 0$ (case 2). Assuming that $\mathbf{f}_{\dot{m}^{\prime\prime}}$ behaves similar to $\mathbf{f}_{\sigma}$, it means the discretization of $\mathbf{f}_{NC}$ is first-order at best. This is a result of the stability (i.e., source term smoothness) trade-off between a CSF-like discretization and the approach described in Section 3.3 using staggered volume fractions to obtain the interfacial terms in $\mathbf{f}_{NC}$.
@@
Despite capturing the correct volume decrease and maintaining a spherical shape, grid convergence of the droplet’s volume is hardly achieved with the plane shift strategy described in Section 2.2. This is a clear consequence of the lack of smoothness in $\mathbf{n}_{\Gamma}$. For instance, Figure 10(b) shows that the method benefits from the smoother $\mathbf{n}_{\Gamma}$ distribution when the spurious currents in $\Omega_{\Gamma}$ are less, such as when $\dot{m}^{\prime\prime}\rightarrow 0$ (case 1) or $\sigma\rightarrow 0$ (case 2). However, since grid refinement demands a larger number of time steps to reach the same $t^{*}=0.08$, the accumulation of geometrical errors during plane shifts prevents a clear converging trend. This is supported by the addition of case 3 in Figure 10(b), which corresponds to case 3 (or case 2) but without solving the momentum equation in the algorithm. That is, the effects of surface tension are not included and the pressure is only used to impose the Stefan flow. This eliminates the spurious currents caused by surface tension and other CSF-like forces, and results in lower volume errors and better convergence. However, the convergence rate is still lower than first order, highlighting the limitations of plane shifting at high $\dot{m}^{\prime\prime}$. Nonetheless, all mesh resolutions from Table 1 display volume errors around 1% or less, in line with other works using different methodologies to handle phase change in VOF methods (e.g., advection of $C$ with $\mathbf{u}_{\Gamma}$ [11, 35]).
Figure 11: Temporal evolution of droplet sphericity ($\Lambda$) in the static evaporating droplet with mesh L3. Case 3* corresponds to case 3 but without solving the momentum equation.
Geometry errors are also measured in terms of the sphericity of the vaporizing droplet, $\Lambda=[\pi^{1/3}(6V)^{2/3}]/A$ where $V$ is the droplet volume and $A$ its surface area calculated as the sum of the area of all the local interface reconstruction planes. The temporal evolution of $\Lambda$ is shown in Figure 11 for all cases with mesh L3, including case 3. The data extends beyond the time reported in Figure 10 and clearly shows the enhanced accumulation of geometrical errors as $\dot{m}^{\prime\prime}$ increases (case 1 vs. case 3). Without the restoring effect of surface tension (case 2), the geometrical errors may degrade the spherical shape completely. Further proof that this is a result of the spurious currents generated around the interface and not the plane shifting strategy is given by case 3, which preserves the spherical shape of the vaporizing droplet better over time. Given these results, a clear connection between the deterioration of $\Lambda$ and the volume convergence errors is observed.
@@
Note that this validation test imposes a restrictive CFL condition based on a maximum allowed interface shift due to phase change in a given time step of $(\Delta\text{d}_{\text{max}}/\Delta x)=0.001$. However, no significant improvement in the volume errors is observed by varying $(\Delta\text{d}_{\text{max}}/\Delta x)$ or $\theta_{3}$ from $0.04$ to $0.001$ (see C for the definition of $\theta_{3}$). Figure 10(c) shows the volume errors for case 3 and, similar to Figure 10(b), grid convergence is not achieved. Except for mesh L1, volume errors are almost identical regardless of $\Delta\text{d}_{\text{max}}$. Thus, a maximum plane shift around $(\Delta\text{d}_{\text{max}}/\Delta x)=0.01$ is more reasonable to avoid reducing the time step significantly if its calculation is limited by the evaporation of the interface. Beyond the reported trends, reducing the time step delays the onset of numerical instabilities because of two main reasons: (a) smaller plane shifts in the phase change correction of $C$ deteriorate less the geometry; and (b) the errors in the temporal extrapolation of $\hat{p}$, $\hat{\mathbf{f}}_{\sigma}$ and $\hat{\mathbf{f}}_{\dot{m}^{\prime\prime}}$ are reduced. Thus, spurious currents are smaller and geometrical calculations do not degrade as fast.
By pushing the solver to its limits in the context of a static configuration, some limitations of the proposed treatment of phase change in the geometric advection of $C$ and the discretization of the various corrective source terms added in the momentum equation have been highlighted. Nonetheless, the magnitude of $\dot{m}^{\prime\prime}$ is typically smaller, surface tension effects prevail (unless very close to the critical point of the fluid), and dynamic configurations aid in masking the spurious currents.